Two parallel wires are seperated by 0.6 meter and carry anti-parallel current of 7 amp each. What is the resultant field between the wires at 0.4 m from lower wire.
Answer #1

Let us assume current is flowing from left to right in upper wire and opposite in lower one.

Magnetic field due to upper wire at a distance 0.2 m from it,

B_{upper}=\frac{\mu_{0}I}{2\pi a}=\frac{7\mu_{0}}{2\pi \0.2} directed towards the plane of paper.

Magnetic field due to lower wire at a distance 0.4 m from it,

B_{lower}=\frac{7\mu_{0}}{2\pi \0.4} directed towards the plane of paper.

So the total field,

B=B_{upper}+B_{lower}=\frac{21\mu_{0}}{0.8\pi }=\frac{21\times 4\pi\times 10^{-7}}{0.8\pi}=1.05\times 10^{-5} T

Magnetic field due to upper wire at a distance 0.2 m from it,